# Practical trainable temporal post-processor for multi-state quantum measurement

### v2

@@ Abstract
We develop and demonstrate a trainable temporal post-processor (TPP) harnessing a simple but versatile machine learning algorithm to provide optimal processing of quantum measurement data subject to arbitrary noise processes, for the readout of an arbitrary number of quantum states. We demonstrate the TPP on the essential task of qubit state readout, which has historically relied on temporal processing via matched filters in spite of their applicability only for specific noise conditions. Our results show that the TPP can reliably outperform standard filtering approaches under complex readout conditions, such as high power readout. Using simulations of quantum measurement noise sources, we show that this advantage relies on the TPP’s ability to learn optimal linear filters that account for general quantum noise correlations in data, such as those due to quantum jumps, or correlated noise added by a phase-preserving quantum amplifier. We show that the transformation described by the TPP can be expressed via an efficient semi-analytic form, providing a linearly-scaling generalization of matched filtering to an arbitrary number of states under the most general noise conditions of the readout signal emanating from the measurement chain. The TPP can be efficiently, autonomously, and reliably trained on measurement data, and requires only linear operations, making it ideal for FPGA implementations in cQED for real-time processing of measurement data from general quantum systems.
@@ I Introduction
High fidelity quantum measurement is essential for any quantum information processing scheme, from quantum computation to quantum machine learning. However, while measurement optimization has focused on quantum hardware advancements [1, 2, 3], several modern experiments operate in regimes where optimal hardware conditions are difficult to sustain, or - for machine learning with general quantum systems [4, 5, 6, 7, 8] - may not always be known. For example, in the push towards higher qubit readout fidelities with complex multi-qubit processors in circuit QED (cQED), optimization of individual readout resonators becomes increasingly difficult. More importantly, finite qubit coherence means that simply extending the measurement duration is not a viable option to enhance fidelity: faster and hence higher power measurements are needed. However, these readout powers are associated with enhanced qubit transitions, leading to the $T_{1}$ versus $\bar{n}$ problem [9, 10, 11, 12, 13, 14, 15, 15] and excitation to higher states [16, 14, 17] outside the computational subspace. Machine learning with quantum devices operating in unconventional regimes allows for an even broader range of complex dynamics. Quantum measurement data obtained under these conditions cannot be expected to be optimally analyzed using schemes built for more standard readout paradigms [18]. Therefore, a practical approach to extract the maximum information possible from such data is timely.
@@
In this paper, we demonstrate a machine learning scheme to optimally process quantum measurement data for completely general quantum state classification tasks. For the most common such task of single-shot qubit state readout, standard post-processing of measurement records has remained relatively unchanged (with some exceptions [19, 20]): data is filtered using a “matched filter” (MF) constructed from the sample mean of measurement records for two states to be distinguished (for example, states |$\ket{e}$ or |$\ket{g}$ of a qubit). Crucially, the MF thus defined applies only to binary classification, and much more restrictively, is optimal only under asymptotic conditions under which readout signal is subject to Gaussian white (i.e. uncorrelated) noise processes [21]. In many deployments where complex conditions prevail (such as multi-qubit readout) an even simpler and less optimal boxcar filter is employed, due to the ease of its construction. Our approach harnesses machine learning to provide a model-free trainable temporal post-processor (TPP) of quantum measurement data under most general noise conditions, and for an arbitrary number of states of a generic measured quantum system (111source code available at https://zenodo.org/doi/10.5281/zenodo.10020462 for source code). We test our approach by applying it to the experimental readout of distinct qubits across a range of measurement powers. Our results demonstrate that the TPP reliably outperforms the standard MF whenever measured data exhibits nontrivial temporal correlations, including those of a quantum origin. We find an important such regime that has garnered significant recent attention to be that of high-power readout [16, 23, 24, 25, 26]; here, we experimentally show that the TPP can provide a reduction in errors by a factor of several in certain cases. Furthermore, the TPP achieves this improvement while requiring only linear weights applied to quantum measurement data (see Fig. 1): this makes it compatible with FPGA implementations for real-time hardware processing, and exacts a lower training cost [27, 28] than neural network-based machine learning schemes [20, 29, 30].
@@ II.2 Quantum noise in dispersive qubit readout
$I^{(\sigma)}(t_{i})$ $=\sqrt{\kappa}\langle\hat{X}^{(\sigma)}(t_{i})\rangle+\xi_{I}(t_{i})+\xi_{I}^{\rm qm}(t_{i})+\xi_{I}^{\rm cl}(t_{i}),$ (8a)
$Q^{(\sigma)}(t_{i})$ $=\sqrt{\kappa}\langle\hat{P}^{(\sigma)}(t_{i})\rangle+\xi_{Q}(t_{i})+\xi_{Q}^{\rm qm}(t_{i})+\xi_{Q}^{\rm cl}(t_{i}).$ (8b)
@@
More importantly, $\xi_{I/Q}^{\rm qm}(t_{i})$ describes quantum noise contributions to measurement records, whose origin is intrinsically tied to the nature of quantum measurement. The measurement of a quantum system imposes an evolution of its state, so that a given measurement affects the outcome of subsequent measurements. This effect is described via a measurement-conditioned stochastic quantum state $\hat{\rho}_{c}$ (referred to as a quantum trajectory), which is distinct from the unconditional quantum state $\hat{\rho}$ formally obtained when ensemble-averaging over repeated measurements. Consequently, for any given measurement instance, observables such as the conditional quadrature expectation $\langle\hat{X}^{(\sigma)}(t_{i})\rangle_{c}={\rm Tr}\{\hat{X}\hat{\rho}_{c}^{(\sigma)}(t_{i})\}$ under heterodyne monitoring can deviate from the unconditional ensemble average $\langle\hat{X}^{(\sigma)}(t_{i})\rangle$; this difference, given by $\xi_{I}^{\rm qm}(t_{i})=\langle\hat{X}^{(\sigma)}(t_{i})\rangle_{c}-\langle\hat{X}^{(\sigma)}(t_{i})\rangle$, manifests as quantum noise. These terms include amplified quantum fluctuations when measuring the output field from a quantum amplifier (see Sec. V.1), or the influence of quantum jumps in the measured cavity field due to transitions of the dispersively coupled qubit (see Sec. V.2). Finally, $\xi_{I/Q}^{\rm cl}(t_{i})$ describe classical noise contributions to measurement records, for example noise added by classical HEMT amplifiers. While the statistics of this noise may take different forms, it is formally distinct from heterodyne measurement noise, as it has no associated stochastic measurement superoperator in Eq. (6).
The objective of the qubit readout task is then to use this noisy single-shot [43] temporal measurement data to obtain an estimated class label $\sigma^{\rm est}$ that is ideally equal to the true class label $\sigma$. Within the TPP framework, then, $N_{\rm O}=2$ and $\bm{\vec{x}}^{(\sigma)}=\begin{pmatrix}\vec{I}^{(\sigma)}\\ \vec{Q}^{(\sigma)}\end{pmatrix}$, where $\vec{I}_{i}=I(t_{i})$. The noise $\bm{\vec{\zeta}}$ in Eq. (3) then contains the terms $\xi,\xi^{\rm qm},\xi^{\rm cl}$. However, before describing TPP results, we first briefly review standard approaches to qubit state classification.
@@ III TPP learning as optimal filtering: Generalized matched filters
When compared against Eq. (10), the interpretation of $\bm{\vec{f}}_{k}$ becomes clear: this set of weights can be viewed as a temporal filter applied to the data $\bm{\vec{x}}$. TPP based classification can therefore be interpreted as the application of $C$ filters (one for each $k$) to obtain the estimated label $\sigma^{\rm est}$. The optimal $\mathbf{W}^{\rm opt}$ therefore defines the optimal filters that enable this estimation with minimal error. The use of $C$ optimal filters for a $C$-state classification task indicates the linear scaling of the TPP approach with the complexity of the task.
@@ III.1 TPP performance under Gaussian white noise in comparison to standard FGDA
Figure 2: Multi-state ($C=3$) classification performance of TPP versus FGDA under Gaussian white noise conditions. We consider dispersive qubit readout to distinguish states $p\in\{e,g,f\}$ as a function of measurement power. For a transmon $\chi_{p}/\kappa\in\{-\chi,\chi,-3\chi\}$, $\chi/\kappa=0.195$, and $\kappa/2\pi=1.54\leavevmode\nobreak\ {\rm MHz}$. More opaque markers indicate stronger measurement tone amplitudes. Inset shows induced dispersive shifts for each state (not to scale). Standard FGDA is performed using one of three MFs corresponding to each distinct state pair, as well as a boxcar filter. TPP filters are also followed by a Gaussian discriminator for an equivalent comparison. Only one of the binary MFs allows the FGDA to approach the TPP, while all other chosen filters yield a worse performance.
@@ IV.1 Experimental Results
Figure 3: Classification performance of TPP versus FGDA for readout of real qubits. (a) Parameters of various dispersive qubit-cavity systems used for gathering readout data. Coherence measurements are subject to $10\%$ variation over time. (b) Representative qubit readout histograms under boxcar filtering as a function of measurement signal amplitude. (c) Readout data for three dispersive qubit-cavity systems is analyzed and the resulting classification infidelities for binary ($C=2$) state classification are plotted against each other. The dashed line marks $1-\mathcal{F}_{\rm FGDA}=1-\mathcal{F}_{\rm TPP{}}$. For datasets with variable shading of markers (red and black), more opaque markers indicate stronger measurement tone amplitudes, with corresponding resonator photon number $\bar{n}$ indicated via colorbars. Inset: Percentage fewer errors $\mathcal{E}$ computed for indicated datasets with increasing input signal amplitude.
@@
In our experiments, we perform readout across this domain using two different qubits. For Qubit A, we simultaneously vary both pulse amplitude and pulse duration ($\mathcal{T}_{\rm off}-\mathcal{T}_{\rm on}$), the latter from 300 ns to 1150 ns, to together obtain roughly 9$\pm$3 to 18$\pm$5 photons in the cavity in the steady state. For Qubit B’s phase-preserving dataset, measurement pulse durations vary independently from 500 ns to 900 ns, and measurement amplitudes are adjusted to drive roughly 44 to 363 photons in the cavity in the steady state; the significantly larger photon number is tolerated due to the low Qubit B $\chi/\kappa$. At the lowest pulse duration and amplitude, this corresponds to just enough discriminating power to separate the measured distributions for the two states by approximately their width in a boxcar-filtered IQ plane (namely, without the use of an empirical MF). An example of the individual readout histograms for qubits initialized in states $p\in\{e,g\}$ at this lowest measurement tone power is shown in Fig. 3(b). Qubit B’s phase-sensitive dataset was recorded with a pulse time of 800ns with a shaped pulse to shorten the effect of the cavity ring-up time similar to [58].
@@
The other two datasets are obtained for readout under varying measurement powers. The depth of shading of the markers indicates the strength of measurement drives: the more opaque the marker, the stronger the measurement power. We first note that the classification fidelity does not uniformly increase with signal amplitude in experiment; this is in contrast to the simplified theoretical simulations of Sec. III.1, and is expected due to the aforementioned dynamical effects exhibited in real qubit readout at higher readout powers (neglected in Fig. 2).
For weaker measurement powers, we see that the TPP and the FGDA are once again comparable. However, a very clear trend emerges: for stronger measurement powers - where measurement dynamics become much more complex as demonstrated in Fig. 3(b) - the TPP generally outperforms the FGDA. To more precisely quantify the difference in performance between the TPP and FGDA, we introduce the metric $\mathcal{E}$:
@@
Figure 4: Adaptation of TPP-learned filters with increasing measurement tone amplitude and evolving noise conditions. Black curves are normalized TPP filters under the white noise assumption; for binary state classification, these are identical to standard matched filters. Gray curves are general TPP filters with no assumptions on noise statistics. (a) Filter $\vec{f}_{1}$ for binary ($C=2$) classification, and (b) filters $\vec{f}_{1,3}$ for $C=3$ state classification. In both cases, at weaker amplitudes the general TPP filter closely matches the TPP filter assuming white noise. However, for stronger measurement amplitudes, a marked difference between the white noise TPP filter and the general TPP filter is observed.
@@ IV.2 Adaptation of TPP-learned filters under strong measurement tones
Clearly, the precise form of filters learned by the TPP to outperform white noise filters must be influenced by some physical phenomena that arise at strong measurement powers. However, the TPP is not provided with any physical description for such phenomena, which is in fact part of its model-free appeal. What then, is the mechanism through which the TPP can learn about such phenomena to compute optimal filters? The answer lies in Eq. (4): TPP-learned filters are sensitive to noise correlations in data via $\mathbf{V}$. Using simulations of measurement chains where the noise structure of quantum measurement data can be precisely controlled, we show that the noise structure can strongly deviate from white noise conditions under practical settings. Crucially, the TPP can adapt to these changes where the MF cannot.

### v3

# Practical Trainable Temporal Postprocessor for Multistate Quantum Measurement

@@ Abstract
We develop and demonstrate a trainable temporal post-processor (TPP) harnessing a simple but versatile machine learning algorithm to provide optimal processing of quantum measurement data subject to arbitrary noise processes, for the readout of an arbitrary number of quantum states. We demonstrate the TPP on the essential task of qubit state readout, which has historically relied on temporal processing via matched filters in spite of their applicability only for specific noise conditions. Our results show that the TPP can reliably outperform standard filtering approaches under complex readout conditions, such as high power readout. Using simulations of quantum measurement noise sources, we show that this advantage relies on the TPP’s ability to learn optimal linear filters that account for general quantum noise correlations in data, such as those due to quantum jumps, or correlated noise added by a phase-preserving quantum amplifier. Furthermore, we derive an exact analytic form for the optimal TPP weights: this positions the TPP as a linearly-scaling generalization of matched filtering, valid for an arbitrary number of states under the most general readout noise conditions, all while preserving a training complexity that is essentially negligible in comparison to that of training neural networks for processing temporal quantum measurement data. The TPP can be autonomously and reliably trained on measurement data and requires only linear operations, making it ideal for FPGA implementations in cQED for real-time processing of measurement data from general quantum systems.
@@ I Introduction
High fidelity quantum measurement is essential for any quantum information processing scheme, from quantum computation to quantum machine learning. However, while measurement optimization has focused on quantum hardware advancements [1, 2, 3], several modern experiments operate in regimes where optimal hardware conditions are difficult to sustain, or - for machine learning with general quantum systems [4, 5, 6, 7, 8] - may not always be known. For example, in the push towards higher qubit readout fidelities with complex multi-qubit processors in circuit QED (cQED), optimization of individual readout resonators becomes increasingly difficult. More importantly, finite qubit coherence means that simply extending the measurement duration is not a viable option to enhance fidelity: faster and hence higher power measurements are needed. However, these readout powers are associated with enhanced qubit transitions, leading to the $T_{1}$ versus $\bar{n}$ problem [9, 10, 11, 12, 13, 14, 15] and excitation to higher states [16, 14, 17] outside the computational subspace. Machine learning with quantum devices operating in unconventional regimes allows for an even broader range of complex dynamics. Quantum measurement data obtained under these conditions cannot be expected to be optimally analyzed using schemes built for more standard readout paradigms [18]. Therefore, a practical approach to extract the maximum information possible from such data is timely.
@@
In this paper, we demonstrate a machine learning scheme to optimally process quantum measurement data for completely general quantum state classification tasks. For the most common such task of single-shot qubit state readout, standard post-processing of measurement records has remained relatively unchanged (with some exceptions [19, 20]): data is filtered using a “matched filter” (MF) constructed from the sample mean of measurement records for two states to be distinguished (for example, states |$\ket{e}$ or |$\ket{g}$ of a qubit). Crucially, the MF thus defined applies only to binary classification, and much more restrictively, is optimal only for idealized conditions under which the readout signal is subject to Gaussian white (i.e. uncorrelated) noise processes [21]. In many deployments where complex conditions prevail (such as multi-qubit readout) an even simpler and less optimal boxcar filter is employed, due to the ease of its construction. Our approach harnesses machine learning to provide a model-free trainable temporal post-processor (TPP) of quantum measurement data under the most general noise conditions, and for an arbitrary number of states of a generic measured quantum system (111source code available at https://zenodo.org/doi/10.5281/zenodo.10020462 for source code). We test our approach by applying it to the experimental readout of distinct qubits across a range of measurement powers. Our results demonstrate that the TPP reliably outperforms the standard MF whenever measured data exhibits nontrivial temporal correlations, including those of a quantum origin. We find an important such regime that has garnered significant recent attention to be that of high-power readout [16, 23, 24, 25, 26]; here, we experimentally show that the TPP can provide a reduction in errors by up to 30% in certain cases. Furthermore, the TPP achieves this improvement while requiring only linear weights applied to quantum measurement data (see Fig. 1): this makes it compatible with FPGA implementations for real-time hardware processing, and exacts a lower training cost [27, 28] than neural network-based machine learning schemes [20, 29, 30].
@@ II.2 Quantum noise in dispersive qubit readout
$I^{(\sigma)}(t_{i})$ $=\sqrt{\kappa}\left[\langle\hat{X}^{(\sigma)}(t_{i})\rangle+\xi_{I}^{\rm qm}(t_{i})\right]+\xi_{I}(t_{i})+\xi_{I}^{\rm cl}(t_{i}),$ (8a)
$Q^{(\sigma)}(t_{i})$ $=\sqrt{\kappa}\left[\langle\hat{P}^{(\sigma)}(t_{i})\rangle+\xi_{Q}^{\rm qm}(t_{i})\right]+\xi_{Q}(t_{i})+\xi_{Q}^{\rm cl}(t_{i}).$ (8b)
@@
More importantly, $\xi_{I}^{\rm qm}(t_{i}),\xi_{Q}^{\rm qm}(t_{i})$ describes quantum noise contributions to measurement records, whose origin is intrinsically tied to the nature of quantum measurement. The measurement of a quantum system imposes an evolution of its state, so that a given measurement affects the outcome of subsequent measurements. This effect is described via a measurement-conditioned stochastic quantum state $\hat{\rho}_{c}$ (referred to as a quantum trajectory), which is distinct from the unconditional quantum state $\hat{\rho}$ formally obtained when ensemble-averaging over repeated measurements. Consequently, for any given measurement instance, observables such as the conditional quadrature expectation $\langle\hat{X}^{(\sigma)}(t_{i})\rangle_{c}={\rm Tr}\{\hat{X}\hat{\rho}_{c}^{(\sigma)}(t_{i})\}$ under heterodyne monitoring can deviate from the unconditional ensemble average $\langle\hat{X}^{(\sigma)}(t_{i})\rangle$; this difference, given by $\xi_{I}^{\rm qm}(t_{i})=\langle\hat{X}^{(\sigma)}(t_{i})\rangle_{c}-\langle\hat{X}^{(\sigma)}(t_{i})\rangle$, manifests as quantum noise. These terms include amplified quantum fluctuations when measuring the output field from a quantum amplifier (see Sec. V.1), or the influence of quantum jumps in the measured cavity field due to transitions of the dispersively coupled qubit (see Sec. V.2). Finally, $\xi_{I/Q}^{\rm cl}(t_{i})$ describe classical noise contributions to measurement records, for example noise added by classical HEMT amplifiers. While the statistics of this noise may take different forms, it is formally distinct from heterodyne measurement noise, as it has no associated stochastic measurement superoperator in Eq. (6).
The objective of the qubit readout task is then to use noisy single-shot [43] temporal measurement data to obtain an estimated class label $\sigma^{\rm est}$ that is ideally equal to the true class label $\sigma$. Within the TPP framework, then, $N_{\rm O}=2$ and $\bm{\vec{x}}^{(\sigma)}=\begin{pmatrix}\vec{I}^{(\sigma)}\\ \vec{Q}^{(\sigma)}\end{pmatrix}$, where $\vec{I}_{i}=I(t_{i})$. The noise $\bm{\vec{\zeta}}$ in Eq. (3) then contains the terms $\xi,\xi^{\rm qm},\xi^{\rm cl}$. However, before describing TPP results, we first briefly review standard approaches to qubit state classification.
@@ III TPP learning as optimal filtering: Generalized matched filters
When Eq. (13) is compared against Eq. (10), the interpretation of $\bm{\vec{f}}_{k}$ becomes clear: this set of weights can be viewed as a temporal filter applied to the data $\bm{\vec{x}}$. TPP based classification can therefore be interpreted as the application of $C$ filters (one for each $k$) to obtain the estimated label $\sigma^{\rm est}$. The optimal $\mathbf{W}^{\rm opt}$ therefore defines the optimal filters that enable this estimation with minimal error. The use of $C$ optimal filters for a $C$-state classification task indicates the linear scaling of the TPP approach with the complexity of the task.
@@ III.1 TPP performance under Gaussian white noise in comparison to standard FGDA
Figure 2: Multi-state ($C=3$) classification performance of TPP versus FGDA under Gaussian white noise conditions. We consider dispersive qubit readout to distinguish states $p\in\{e,g,f\}$ as a function of measurement power. For a transmon $\chi_{p}/\kappa\in\{-\chi,\chi,-3\chi\}$, $\chi/\kappa=0.195$, and $\kappa/2\pi=1.54\leavevmode\nobreak\ {\rm MHz}$. More opaque markers indicate higher measurement tone amplitudes. Inset shows induced dispersive shifts for each state (not to scale). Standard FGDA is performed using one of three MFs corresponding to each distinct state pair, as well as a boxcar filter. TPP filters are also followed by a Gaussian discriminator for an equivalent comparison. Only one of the binary MFs allows the FGDA to approach the TPP, while all other chosen filters yield a worse performance.
@@ IV.1 Experimental Results
Figure 3: Classification performance of TPP versus FGDA for readout of real qubits. (a) Parameters of various dispersive qubit-cavity systems used for gathering readout data. Coherence measurements are subject to $10\%$ variation over time. (b) Representative qubit readout histograms under boxcar filtering as a function of measurement signal amplitude. (c) Readout data for three dispersive qubit-cavity systems is analyzed and the resulting classification infidelities for binary ($C=2$) state classification are plotted against each other. The dashed line marks $1-\mathcal{F}_{\rm FGDA}=1-\mathcal{F}_{\rm TPP{}}$. For datasets with variable shading of markers (red and black), more opaque markers indicate higher measurement tone amplitudes, with corresponding resonator photon number $\bar{n}$ indicated via colorbars. Inset: Percentage fewer errors $\mathcal{E}$ computed for indicated datasets with increasing input signal amplitude.
@@
In our experiments, we perform readout across this domain using two different qubits. For Qubit A, we simultaneously vary both pulse amplitude and pulse duration ($\mathcal{T}_{\rm off}-\mathcal{T}_{\rm on}$), the latter from 300 ns to 1150 ns, to together obtain roughly 9$\pm$3 to 18$\pm$5 photons in the cavity in the steady state. For Qubit B’s phase-preserving dataset, measurement pulse durations vary independently from 500 ns to 900 ns, and measurement amplitudes are adjusted to drive roughly 44$\pm$5 to 363$\pm$40 photons in the cavity in the steady state; the significantly larger photon number is tolerated due to the low Qubit B $\chi/\kappa$. At the lowest pulse duration and amplitude, this corresponds to just enough discriminating power to separate the measured distributions for the two states by approximately their width in a boxcar-filtered IQ plane (namely, without the use of an empirical MF). An example of the individual readout histograms for qubits initialized in states $p\in\{e,g\}$ at this lowest measurement tone power is shown in Fig. 3(b). Qubit B’s phase-sensitive dataset was recorded with a pulse time of 800ns with a shaped pulse to shorten the effect of the cavity ring-up time similar to [58].
@@
The other two datasets are obtained for readout under varying measurement powers. The depth of shading of the markers indicates the strength of measurement drives: the more opaque the marker, the higher the measurement power. We first note that the classification fidelity does not uniformly increase with signal amplitude in experiment; this is in contrast to the simplified theoretical simulations of Sec. III.1, and is expected due to the aforementioned dynamical effects exhibited in real qubit readout at higher readout powers (neglected in Fig. 2).
For weaker measurement powers, we see that the TPP and the FGDA are once again comparable. However, a very clear trend emerges: for higher measurement powers - where measurement dynamics become much more complex as demonstrated in Fig. 3(b) - the TPP generally outperforms the FGDA. To more precisely quantify the difference in performance between the TPP and FGDA, we introduce the metric $\mathcal{E}$:
@@
Figure 4: Adaptation of TPP-learned filters with increasing measurement tone amplitude and evolving noise conditions. Black curves are normalized TPP filters under the white noise assumption; for binary state classification, these are identical to standard matched filters. Gray curves are general TPP filters with no assumptions on noise statistics. (a) Filter $\vec{f}_{1}$ for binary ($C=2$) classification, and (b) filters $\vec{f}_{1,3}$ for $C=3$ state classification. In both cases, at weaker amplitudes the general TPP filter closely matches the TPP filter assuming white noise. However, for higher measurement amplitudes, a marked difference between the white noise TPP filter and the general TPP filter is observed.
@@ IV.2 Adaptation of TPP-learned filters under strong measurement tones
Clearly, the precise form of filters learned by the TPP to outperform white noise filters must be influenced by some physical phenomena that arise at strong measurement powers. However, the TPP is not provided with any physical description for such phenomena, which is in fact part of its model-free appeal. What then, is the mechanism through which the TPP can learn about such phenomena to compute optimal filters? The answer lies explicitly in Eq. (14): TPP-learned filters are sensitive to noise correlations in data via $\mathbf{V}$. Using simulations of measurement chains where the noise structure of quantum measurement data can be precisely controlled, we show that the noise structure can strongly deviate from white noise conditions under practical settings. Crucially, the TPP can adapt to these changes whereas the MF cannot.